# Early stages of collective cell invasion: Biomechanics

### v1

@@ Abstract
The early stages of the collective invasion may occur by single mesenchymal cells or hybrid epithelial-mesenchymal cell groups that detach from cancerous tissue. Tumors may also emit invading protrusions of epithelial cells, which could be led (or not) by a basal cell. Here we devise a fractional step cellular Potts model comprising passive and active cells able to describe these different types of collective invasion before cells start proliferating. Durotaxis and active forces have different symmetry properties and are included in different half steps of the fractional step method. Compared with a single step method, fractional step produces more realistic cellular invasion scenarios with little extra computational effort. Biochemical mechanisms that determine how cells acquire their different phenotypes and cellular proliferation will be incorporated to the model in future publications.
1. Introduction
@@
(ii) Collective invasion may occur in different ways. Hybrid epithelial/mesenchymal (E/M) cells are related to different stages of the EMT pas18 ; bra18 , and migrate as small clusters. M and E/M cells at different stages of the EMT are found in a variety of tumors pas18 . (iii) Cancer tissue forms protrusions comprising a number of E cells that are alike, in an instability similar to finger formation in wound healing hak17 , or (iv) the protrusions are led by basal cells. Basal tumor cells have E phenotype and lead invasion without undergoing the EMT cho14 . Biochemical cues, EMT and cellular Notch signaling may trigger the invasion and they are intimately connected to biomechanical cellular motion and interaction boc18 ; vil21 ; vil22 ; muk22 .
Here we focus on the mechanical aspects of cellular motion and the early stages of invasion before cells proliferate. Our work provides a proof-of-concept biomechanics framework for invasion prior to cell proliferation, distinguishing it from biochemical EMT or signaling models. Starting from a cellular Potts model (CPM) Graner ; Glazier that includes durotaxis vanOers , we add novel active forces that affect differently to epithelial, mesenchymal and hybrid cells: each cell type has a specific local stiffness threshold that characterizes its response to strain gradients. Thus, the EMT acts as a transition between different mechanical regimes: E cells are static and very adhesive, M cells are active, less adhesive and freer to move, whereas E/M hybrid cells combine both behaviors.
@@
When implementing the mechanism involving interconnections among EMT, cancer stem cells and Notch-Jagged signaling boc18 , it is important to characterize the mechanics of different cell phenotypes. Bocci et al have modeled EMT, cancer stem cells and Notch signaling by sets of interconnected ordinary differential equations (ODEs) boc18 . They have studied phenotypes associated to each of the three cycles and explained how treatments that inhibit some variable perform boc18 . Distinguishing phenotypes and assigning them mechanical properties is of paramount importance when using a realistic computational tool such as the cellular Potts model (CPM) to characterize cellular motion Graner ; Glazier ; vanOers . For example, in early stage angiogenesis, the phenotype tip (leader), stalk (follower) or hybrid tip/stalk (leader/follower) cell depends on biochemical factors such as the content of VEGF and Delta protein in the cells and is determined by the ODEs of Notch signaling dynamics boa15 ; jol15 ; boa16 . When using the CPM, the parameters measuring the phenotype sensitivity to mechanical and chemical cues as well as the ability to proliferate determine the behavior of the cells veg20 ; veg21 . It is crucial to appropriately select the values of the different CPM parameters and to check the effects that changing them have on cell behavior veg20 . Once this is known, appropriate parameter values and qualitative properties need to be assigned to cells according to their phenotype: sensitivity to chemotaxis and proliferation of tip, stalk and hybrid tip/stalk cells in the case of angiogenesis veg20 .
@@
Collective cell invasion depends on the ECM structure. Its role in cell migration and morphogenesis is discussed in Ref. cro24 . The interstitial ECM fills spaces between organs and it has fibrous proteins and proteoglycans that form 3D arrangements. The basement membrane ECM forms 2D sheet-like ECM that line organ boundaries cro24 . The influence of a fibrous, cross-linked ECM (modeled by bead-springs) on cell motion modeled by a CPM is discussed in Ref. tsi23 whereas the emergence of different cell motility types from cell-matrix adhesion dynamics via a modified CPM is shown in Ref. vst22 . For collective invasion of cancer cells, the relation between phenotypes within the EMT, Notch signaling and cancer stem cell cycles boc18 , and the mechanical part of the CPM has not been established. Recently, Hirway et al have incorporated EMT to the CPM of Ref. vanOers by means of a set of ODEs for one type of cells and for the ECM hir21 and for several cell types and for the ECM hir24 . These ODEs include source terms dependent on junctional forces that are coupled to the equations of elasticity comprising the mechanical part of the model hir21 . They determine the E, M and E/M phenotype of the cells, and each phenotype has different adhesion and proliferation properties. The resulting model has been used to describe tissue formation and a wound healing (scratch) assay in which the cells of a tissue close a space that has been opened in their midst hir21 . In these simulations, the cells have to proliferate to complete the tissue or to close the opening. However, openings in experiments with epithelial tissue often close without cellular proliferation, either in wound healing assays pet10 or in antagonistic migration assays moi19 .
@@
Different experiments have shown different phases in migrating clusters of malignant cells that can be modeled using ideas from active matter cop18 . Experiments, theory and numerical simulations have shown the existence of chiral currents in confined monolayers of highly active spindle-shaped human fibrosarcoma cells with sources of active forces at the interfaces yas22 . Long range force transmission of mechanical origin fosters phenomena such as waves in antagonistic migration assays rod17 , stress fibres mechanics oak17 and collective cell durotaxis sun16 . Is it possible to include active forces in the CPM? Rens and Edelstein-Keshet have considered forces due to cell-cell adhesion and also active forces due to a two-state signaling protein inside cells: they modify the CPM energy to incorporate the solutions of reaction-diffusion equations for the concentrations of active/inactive protein states ren19 . Their results include measured polarization of the cells and retraction or expansion of the cell boundary ren19 . Including pseudopods in the CPM can be achieved by defining a persistence variable related to actin and adding appropriate terms to the CPM energy bur22 .
@@
In this paper, we consider a circular aggregate comprising epithelial, mesenchymal and hybrid E/M cells and study their migration and invasion at times shorter than that of cellular division. Mesenchymal and hybrid E/M cells are subject to active forces that do not affect epithelial cells. Cells with different phenotypes have different adhesion parameters. Biochemical cues that include coupled EMT, cancer stem cell and Notch signaling dynamics will be incorporated in a subsequent paper. Our results show that mesenchymal and hybrid cells placed near the edges of the aggregate escape toward an attraction point as single cells and small aggregates, respectively. Hybrid cells are more effective in their migration. Our work is a proof of concept that a combination of cellular phenotypes and mechanical properties can model early individual and collective invasion of tissues with possible applications for tumors and cancer metastasis.
2. Dynamics of Cells on a Substrate
@@
For the biomechanics of invading cells, we will consider durotaxis, the propensity of cells to migrate up gradients of substrate rigidity vanOers ; lo00 , and active forces that direct cells to migrate to other regions. Following van Oers et al vanOers , we assume that a strained ECM is stiffer along the strain orientation than perpendicular to it (strain stiffening), that the ECM is isotropic and linearly elastic, and that stiffness is an increasing, linear function of the local strain. To mimic the attraction of cells to a particular region, we shall assume that they are pushed towards the target by a repelling linear force.
2.1. Influence of Substrate on Cell Dynamics
To study the influence of substrate stiffness on cell behavior, we build upon the computational model introduced in Ref. vanOers to which we add active migration forces.
2.1.1. Durotaxis
@@
Durotaxis is modeled as in Ref. vanOers . We solve the Navier equations for isotropic linear elasticity using the finite element method (FEM) as follows:
$\mathbf{K}\mathbf{u}$ $=$ $\mathbf{f},\quad\mathbf{f}_{i}=\mu\sum_{j}\mathbf{d}_{ij}.$ (1)
@@
Here $\mathbf{K}$ is the stiffness matrix, $\mathbf{u}$ the displacement vector at each node, and $\mathbf{f}$ the force vector. The force $\mathbf{f}_{i}$ at node $i$ is the tension per unit length, $\mu$, multiplied by the sum of distances $\mathbf{d}_{ij}=\mathbf{n}_{j}-\mathbf{n}_{i}$ between node $i$ and its neighboring nodes $j$, where $\mathbf{n}_{i}$ and $\mathbf{n}_{j}$ are their respective positions. The force $\mathbf{f}_{i}$ mimics the cell-shape dependent contractile forces exerted by the cells onto the ECM, it does not derive from a potential energy, and it gives experimentally plausible predictions for fibroblasts, endothelial cells, and keratocytes lem10 .
@@
In the present CPM, the pixels $\mathbf{x}$ of a square grid (or in a cube in 3D) have labels identifying them as part of a cell or the ECM. After a random pixel change $\mathbf{x}\to\mathbf{x^{\prime}}$, the CPM uses the Metropolis algorithm to update an energy $H$ defined on the pixels. The energy comprises a term enforcing cell volume, a contact term modeling cell-cell and cell-substrate adhesion and a term monitoring changes in strain. The probability of accepting a change $\Delta H$ in energy is 1 if $\Delta H<0$ and $e^{-\Delta H/T}$ otherwise:
@@
$\Delta H=\Delta H_{volume}+\Delta H_{contact}+\Delta H_{strain}.$ (3)
We have
$H_{volume}$ $=$ $\sum_{\sigma\in cells}\lambda\left(\frac{a(\sigma)-A(\sigma)}{A(\sigma)}\right)^{2},$ (4)
$H_{contact}$ $=$ $\sum_{\mathbf{x},\mathbf{x}^{\prime}}J(\sigma(\mathbf{x}),\sigma(\mathbf{x}^{\prime}))\left(1-\delta(\sigma(\mathbf{x}),\sigma(\mathbf{x}^{\prime}))\right).$ (5)
Here $a(\sigma)$ is the actual cell volume, $A(\sigma)$ the target (resting) volume, $\sigma$ an integer representing the label of each cell and substrate, and $\lambda$ is the strength of the volume constraint. $J(\sigma(\mathbf{x}),\sigma(\mathbf{x}^{\prime}))\geq 0$ represents the adhesion strength between neighboring pixels $\mathbf{x}$ and $\mathbf{x}^{\prime}$, while $\delta$ is the Kronecker delta. Adhesion values for cell-cell interactions are larger than those for cell-substrate (e.g., $J_{cc}\approx 2J_{cs}=2.5$).
@@
$\Delta H_{strain}$ $=$ $-g(\mathbf{x},\mathbf{x}^{\prime})\lambda_{\text{duro}}\left[h(E(\epsilon_{1}))(\mathbf{v}_{1}\cdot\mathbf{v}_{m})^{2}+h(E(\epsilon_{2}))(\mathbf{v}_{2}\cdot\mathbf{v}_{m})^{2}\right]\!.$ (6)
Here $g(\mathbf{x},\mathbf{x}^{\prime})$ distinguishes extension (+1+1) and retraction (1-1), $\lambda_{\text{duro}}$ controls the strain contribution, and $\mathbf{v}_{m}$ is a unit vector in the copy direction $\mathbf{x}-\mathbf{x}^{\prime}$. The eigenvalues $\epsilon_{1}$ and $\epsilon_{2}$ of the strain tensor are the principal strains, and the corresponding eigenvectors, $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$, provide the strain orientation. Writing now the strain tensor as a column vector, we have
@@
where $\epsilon_{\text{st}}$ adjusts stiffening. The sigmoidal function $h(E)$ regulates focal adhesion maturation based on threshold stiffness $E_{\theta}$:
$h(E)$ $=$ $\frac{1}{1+e^{-\beta(E(\epsilon)-E_{\theta})}},$ (9)
where $\beta$ is the steepness of the sigmoid. In extension, $g=1$ in Eq. (6), we use the value $E_{\theta}$ of the target pixel $\mathbf{x}^{\prime}$, whereas we use the value $E_{\theta}$ of the source pixel $\mathbf{x}$ in retraction, g=1g=-1 in Eq. (6). This symmetry extension-retraction keeps the cohesion of the cells. Notice that the force vector in Eq. (1) changes the energy in Eq. (6) through its influence on the strain tensor that also modifies the substrate stiffness. The strain-driven energy change of Eq. (3) entering the Monte Carlo step is the effective force that moves the cells.
The parameter values we use agree with the methodology outlined in vanOers . The two-dimensional computational domain (substrate) is defined by its size in pixels, with $NV_{x}$ and $NV_{y}$ representing the number of pixels along each side. Each pixel corresponds to a physical size of $dx=2.5\mu m$ resulting in domain dimensions $Lx=NV_{x}\cdot dx$ and $Ly=NV_{y}\cdot dx$. In this section, we set $NV_{x}=NV_{y}=300$, which gives a domain size of $\mum ×\times 750μ750\mum$. The parameter values listed in Table 1 are compatible with those in Ref. vanOers and capture the qualitative behavior of the system while ensuring numerical stability and consistency with experimental observations. We have included a reference adhesion parameter $J_{\text{ref}}$ indicative of the adhesion moduli for the different cell phenotypes.

### v2

@@ Abstract
The early stages of the collective invasion may occur by single mesenchymal cells or hybrid epithelial-mesenchymal cell groups that detach from cancerous tissue. Tumors may also emit invading protrusions of epithelial cells, which could be led (or not) by a basal cell. Here we devise a novel fractional step cellular Potts model comprising passive and active cells able to describe these different types of collective invasion before cells start proliferating. Cells moving toward stiffness gradients (durotaxis) and active forces pulling them away from the tumor have different symmetry properties under cellular extension and retraction that sometimes hamper collective invasion when put together. Thus, these forces are included in different half steps of the fractional step method. Compared with a single step method, fractional step produces more realistic collective invasion scenarios with little extra computational effort. Biochemical mechanisms that determine how cells acquire their different phenotypes and cellular proliferation will be incorporated to the model in future publications.
I Introduction
@@
(ii) Collective invasion may occur in different ways. Hybrid epithelial/mesenchymal (E/M) cells are related to different stages of the EMT pas18 ; bra18 , and migrate as small clusters. M and E/M cells at different stages of the EMT are found in a variety of tumors pas18 .
(iii) Cancer tissue forms protrusions comprising a number of E cells that are alike, in an instability similar to finger formation in wound healing hak17 , or
(iv) The protrusions are led by basal cells. Basal tumor cells have E phenotype and lead invasion without undergoing the EMT cho14 .
@@
Biochemical cues, EMT and cellular Notch signaling may trigger the invasion and they are intimately connected to biomechanical cellular motion and interaction boc18 ; vil21 ; vil22 ; muk22 . Bocci et al have modeled EMT, cancer stem cells and Notch signaling by sets of coupled ordinary differential equations (ODEs) boc18 . They have studied phenotypes associated to each of the three cycles and explained how treatments inhibiting some variable in the cycles perform boc18 . Distinguishing phenotypes and assigning them mechanical properties is of paramount importance when using a realistic computational tool such as the cellular Potts model (CPM) to characterize cellular motion Graner ; Glazier ; vanOers . For example, in early stage angiogenesis, the phenotype tip (leader), stalk (follower) or hybrid tip/stalk (leader/follower) cell depends on biochemical factors such as the content of VEGF and Delta protein in the cells. These contents are determined by the ODEs of Notch signaling dynamics boa15 ; jol15 ; boa16 . The CPM parameters measuring the phenotype sensitivity to mechanical and chemical cues, as well as the ability to proliferate, determine the behavior of the cells veg20 ; veg21 . It is crucial to appropriately select their values and to check their effects veg20 . Once they are known, appropriate parameter values and qualitative properties need to be assigned to cells according to their phenotype: for angiogenesis, sensitivity to chemotaxis and proliferation of tip, stalk and hybrid tip/stalk cells veg20 .
@@
Collective cell invasion depends on the ECM structure. Its role in cell migration and morphogenesis is discussed in Ref. cro24 . The interstitial ECM fills spaces between organs and it has fibrous proteins and proteoglycans that form 3D arrangements. The basement membrane ECM forms 2D sheet-like ECM that line organ boundaries cro24 . The influence of a fibrous, cross-linked ECM (modeled by bead-springs) on cell motion modeled by a CPM is discussed in Ref. tsi23 . The emergence of different cell motility types from cell-matrix adhesion dynamics via a modified CPM is shown in Ref. vst22 . For collective invasion of cancer cells, the relation between biomechanics and the phenotypes within the EMT, Notch signaling and cancer stem cell cycles boc18 has to be specified. Recently, Hirway et al have incorporated EMT to the CPM of Ref. vanOers , either by means of a set of ODEs for one type of cells and for the ECM hir21 , or for several cell types and for the ECM hir24 . These ODEs include source terms dependent on junctional forces that are coupled to the equations of elasticity comprising the mechanical part of the model hir21 . They determine the E, M and E/M phenotype of the cells, each having different adhesion and proliferation properties. The resulting model describes tissue formation and a wound healing (scratch) assay in which the cells of a tissue close a space that has opened in their midst hir21 . In these simulations, the cells have to proliferate to complete the tissue or to close the opening. However, experiments in wound healing assays pet10 or in antagonistic migration assays moi19 show that openings between epithelial tissues may close before cells have time to proliferate. Thus, it is important to model early stages of collective invasion before cells proliferatate.
@@
Different experiments have shown different phases in migrating clusters of malignant cells that can be modeled using ideas from active matter cop18 . Experiments, theory and numerical simulations have shown the existence of chiral currents in confined monolayers of highly active spindle-shaped human fibrosarcoma cells with sources of active forces at the interfaces yas22 . Long range force transmission of mechanical origin fosters phenomena such as waves in antagonistic migration assays rod17 , stress fibres mechanics oak17 and collective cell durotaxis sun16 . A CPM with migrating cells subject to active Brownian motion describes ameboid motion in a porous material wur26 . Cellular self-propulsion and inhomogeneous polarized cell-cell adhesion affect collective migration mat17 . Cell-cell adhesion and active forces due to a two-state signaling protein inside cells may modify the CPM energy to incorporate the solutions of reaction-diffusion equations for the concentrations of active/inactive protein states ren19 . Including pseudopods in the CPM can be achieved by defining a persistence variable related to actin and adding appropriate terms to the CPM energy bur22 .
@@
Here we focus on the mechanical aspects of cellular motion and the early stages of invasion before cells proliferate. Our work provides a proof-of-concept biomechanics framework for invasion prior to cell proliferation, distinguishing it from biochemical EMT or signaling models. Starting from a CPM that includes durotaxis vanOers , we add novel active forces that are felt only by mesenchymal and hybrid cells, not by epithelial cells. Each step of a novel fractional step Monte Carlo dynamics is divided into two substeps: traction cells that are the basis of durotaxis intervene in the first substep whereas active forces act on the second substep. Each cell type has a specific local stiffness threshold that characterizes its response to strain gradients. E cells are static and very adhesive, M cells are motile, less adhesive and freer to move, whereas E/M hybrid cells combine both behaviors. M and hybrid E/M cells are subject to active forces that do not affect E cells. We consider a circular aggregate comprising E, M and hybrid E/M cells and study their migration and invasion at times shorter than that of cellular division. Our results show that mesenchymal and hybrid cells placed near the edges of the aggregate escape toward an attraction point as single cells and small aggregates, respectively. Hybrid cells are more effective in their migration. Biochemical cues that include coupled EMT, cancer stem cell and Notch signaling dynamics will be incorporated in a subsequent paper. Our work is a proof of concept that a combination of cellular phenotypes and mechanical properties can model early individual and collective invasion of tissues with possible applications for tumors and cancer metastasis.
The rest of the paper is organized as follows. Section II reviews the CPM with durotaxis. Section III explains how cell adhesion and a threshold of the Young modulus affect cellular dynamics and distinguish passive and motile cells. Section IV introduces active migration forces for motile cells and argues that they need to be considered separately in a fractional step CPM to achieve effective migration. Section V describes the different types of collective cellular invasion and the obtained results are discussed in Section VI. In Appendix A, starting from a circular aggregate of passive and motile cells, we enumerate 21 configurations and parameter values that produce seven different patterns. Appendix B contains flowcharts of single step and fractional step CPMs.
II Dynamics of Cells on a Substrate: Durotaxis
@@
Epithelial cells have a propensity to migrate up gradients of substrate rigidity, which is called durotaxis vanOers ; lo00 . In this section, we review van Oers et al’s CPM with durotaxis, which is calibrated with data from experiments vanOers . Thus, we will assume that a strained ECM is stiffer along the strain orientation than perpendicular to it (strain stiffening), that the ECM is isotropic and linearly elastic, and that stiffness is an increasing, linear function of the local strain. Even for homogeneous ECM, strains due to the traction of the cells generate stiffness gradients by modifying the elastic constants (e.g., the Young modulus).
@@
We solve the Navier equations for isotropic linear elasticity using the finite element method (FEM) as follows:
$\mathbf{K}\mathbf{u}$ $=$ $\mathbf{f},\quad\mathbf{f}_{i}=\mu\sum_{j}\mathbf{d}_{ij},\quad\mathbf{d}_{ij}=\mathbf{n}_{j}-\mathbf{n}_{i}.$ (1)
@@
The FEM represents the substrate as a lattice of finite elements $e$, each corresponding to a square pixel of the CPM having four nodes. In Eq. (1), $\mathbf{K}$ is the stiffness matrix, $\mathbf{u}$ the displacement vector, and $\mathbf{f}$ the force vector. The vector $\mathbf{u}=(u_{x_{1}},u_{y_{1}},\ldots,u_{x_{n}},u_{y_{n}})^{T}$ contains the displacements of all nodes, which are the unknowns that the FEM calculates based on the traction forces $\mathbf{f}$ exerted by the cells onto the material. Force $\mathbf{f}_{i}$ at node $i$ (with position $\mathbf{n}_{i}$) is the tension per unit length, $\mu$, multiplied by the sum of distances $\mathbf{d}_{ij}$ between node $i$ and all other nodes nodes $j$ at $\mathbf{n}_{j}$ within the same cell. The force $\mathbf{f}_{i}$ mimics the cell-shape dependent contractile forces exerted by the cells onto the ECM, it does not derive from a potential energy, and it gives experimentally plausible predictions for fibroblasts, endothelial cells, and keratocytes lem10 .
@@
In the present CPM, the pixels $\mathbf{x}$ of a square grid (or in a cube in 3D) have labels identifying them as part of a cell or the ECM. After a random pixel change in which the label of the source $\mathbf{x}$ is copied to the target $\mathbf{x^{\prime}}$, the CPM uses the Metropolis algorithm to update an energy $H$ defined on the pixels. The energy comprises a term enforcing cell volume, a contact term modeling cell-cell and cell-substrate adhesion and a term monitoring changes in strain. The probability of accepting a change $\Delta H$ in energy is 1 if $\Delta H<0$ and $e^{-\Delta H/T}$ otherwise:
@@
$\Delta H=\Delta H_{\text{volume}}+\Delta H_{\text{contact}}+\Delta H_{\text{duro}},$ (3)
$H_{\text{volume}}=\sum_{\sigma\in cells}\lambda\left(\frac{a(\sigma)-A(\sigma)}{A(\sigma)}\right)^{2},$ (4)
$H_{\text{contact}}=\sum_{\mathbf{x},\mathbf{x}^{\prime}}J(\sigma(\mathbf{x}),\sigma(\mathbf{x}^{\prime}))\left(1-\delta(\sigma(\mathbf{x}),\sigma(\mathbf{x}^{\prime}))\right).$ (5)
Here $a(\sigma)$, $A(\sigma)$, $\sigma$, $\lambda$, and $\delta(.,.)$ are the actual cell volume, target volume, an integer representing the label of each cell and substrate, the strength of the volume constraint, and the Kronecker delta, respectively. $J(\sigma(\mathbf{x}),\sigma(\mathbf{x}^{\prime}))\geq 0$ is the adhesion strength between neighboring pixels $\mathbf{x}$ and $\mathbf{x}^{\prime}$, which is larger for cell-cell interactions than for cell-substrate (e.g., $J_{cc}\approx 2J_{cs}=2.5$). Adding a perimeter constraint to Eq. (4) scarcely influences CPM simulation results veg20 and in its effect can be often subsumed in the adhesion parameter dur19 .
@@
$\Delta H_{\text{duro}}$ $=$ $-g(\mathbf{x},\mathbf{x}^{\prime})\lambda_{\text{duro}}\left[h(E(\epsilon_{1}))(\mathbf{v}_{1}\cdot\mathbf{v}_{m})^{2}+h(E(\epsilon_{2}))(\mathbf{v}_{2}\cdot\mathbf{v}_{m})^{2}\right]\!.$ (6)
Here $g(\mathbf{x},\mathbf{x}^{\prime})$ distinguishes extension (+1+1) and retraction (1-1), $\lambda_{\text{duro}}$ controls the strain contribution, and $\mathbf{v}_{m}$ is a unit vector in the copy direction $\mathbf{x}^{\prime}-\mathbf{x}$. The eigenvalues $\epsilon_{1}$ and $\epsilon_{2}$ of the strain tensor are the principal strains, and the corresponding eigenvectors, $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$, provide the strain orientation. Writing now the strain tensor as a column vector, we have
@@
where $\epsilon_{\text{st}}$ adjusts stiffening (which only occurs for extensions $\epsilon>0$, not compressions $\epsilon<0$). The sigmoidal function $h(E)$ regulates focal adhesion maturation based on the Young modulus threshold (YMT) $E_{\theta}$:
$h(E)$ $=$ $\frac{1}{1+e^{-\beta(E-E_{\theta})}},$ (9)
where $\beta$ is the steepness of the sigmoid. In Eq. (6), we use the strain value of the target pixel $\mathbf{x}^{\prime}$ in extension, and that of the source pixel $\mathbf{x}$ in retraction. This extension-retraction symmetry keeps the cohesion of the cells. The force vector in Eq. (1) changes the energy Eq. (6) through the strain tensor that also modifies the substrate stiffness. The strain-driven energy change entering the Monte Carlo time step (MCTS) is the effective force moving the cells.
The two-dimensional computational domain (substrate) is defined by its size in pixels, with $NV_{x}$ and $NV_{y}$ representing the number of pixels along each side. Each pixel corresponds to a physical size of $dx=2.5\mu m$ resulting in domain dimensions $Lx=NV_{x}\cdot dx$ and $Ly=NV_{y}\cdot dx$. In this section, we set $NV_{x}=NV_{y}=300$, which gives a domain size of $\mum ×\times 750μ750\mum$. The parameter values listed in Table 1 are compatible with those in Ref. vanOers , capture the qualitative behavior of the system and ensure numerical stability and consistency with experimental observations. The reference adhesion parameter $J_{\text{ref}}$ is representative of the adhesion moduli for the different cell phenotypes.
@@
III Young modulus threshold and adhesion in durotaxis
Here we consider the behavior of two types of cells that differ in their YMT and/or adhesion properties under the CPM Eqs. (1)-(9). One of the cells will turn out to be passive and remain epithelial cells sharing line boundaries whereas the others will be motile cells that may even leave the tumor. How do we proceed?